Delay Corp. does not currently pay any dividends. The CFO declared that the firm will pay its first annual dividend of $10 per share in 3 years. From there, the firm will grow its annual dividends at 2% per year. If the fair cost of equity for Delay Corp. is 12% per year, what is the fair share price?
Answer #1

This question requires application of constant growth dividend discount model and basic time value of money function,

According to constant growth dividend discount model,

Div Po r-g Po= Price of Stock Div = Estimat ed Dividends for Next Period rRequired Rate of Return g Growth Rate

Div3 Рэ- rg

10 P2 0.12 0.02

100

Now, based on time value of money function,

FV = PV * (1 + r)n

$100 = PV * (1 + 12%)2

$100 = PV * 1.2544

PV = $79.72